OpenStudy (anonymous):

how does -1/3 relate to -3 A. additive inverse B. distributive inverse C. multiplicative inverse D. opposite inverse

OpenStudy (texaschic101):

multiplication inverse -1/3 * -3/1 = 3/3 = 1 The product of a number and its multiplication inverse is 1

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

no problem :)